# On the Fourier Coefficients of Critical Gaussian Multiplicative Chaos

### v1

@@ 4.3. Dominant contribution
It therefore suffices to prove that the right-hand side is $\ll(\log n)^{-2}$ as $t\to\infty$. The bulk of the work lies in proving the following uniform upper bound for $\mathcal{F}_{n,t}$ whose proof is deferred to Section 4.3.
To state it, we introducing the branching time
${r_{\Delta}}=t\land\big(\log|\Delta|^{-1}\lor r_{n}\big),\quad\text{ for }\Delta\in[e^{-t},\Delta_{n}]$
This should essentially be viewed as $\log|\Delta|^{-1}$ on a first reading. The minimum and maximum taken in the definition only serve to make $r_{\Delta}$ well-behaved near the “boundary” (namely $\Delta\approx e^{-t}$ and $\Delta\approx\Delta_{n}$).
@@
Let $\mathcal{F}_{n,t}$ be defined as in Equation (4.18). Then for large enough $t,n>0$, there exists a constant $C(A)=C>0$ such that $\mathcal{F}_{n,t}(\Delta)$ satisfies the bound
(4.20) $\mathcal{F}_{n,t}(\Delta)\leq C\frac{1}{r_{n}^{2}}\frac{t}{(t-{r_{\Delta}})+1}\bigg(\frac{\log(r_{\Delta}-r_{n})^{5}}{(r_{\Delta}-r_{n})^{2}}\land 1\bigg)\frac{1}{\Delta},\quad\text{ for }\Delta\in[e^{-t},\Delta_{n}]$
uniformly in $\Delta$ in said range, and
(4.21) $\mathcal{F}_{n,t}(\Delta)\leq Cf(n)\frac{e^{t}}{\sqrt{t}},\quad\text{ for }\Delta\in[0,e^{-t})$
for some positive function $f$ of $n$.
@@
For large enough fixed $n>0$, we can use Proposition 4.3 to bound the integral in (4.19), beginning with the range $0,e^{-t}]$ where
[…]
On $e^{-t},\Delta_{n}]$, we use the bound in (4.20) to write
$\int_{e^{-t}}^{\Delta_{n}}\mathcal{F}_{n,t}(\Delta)\mathrm{d}\Delta\ll_{A}\frac{1}{{r_{n}}^{2}}\int_{e^{-t}}^{\Delta_{n}}\frac{t}{(t-{r_{\Delta}})+1}\bigg(\frac{\log(r_{\Delta}-r_{n})^{5}}{(r_{\Delta}-r_{n})^{2}}\land 1\bigg)\frac{1}{\Delta}\mathrm{d}\Delta.$
We first take care of the range where $r_{\Delta}$ is close to $r_{n}$. On $\Delta_{n}/e^{3},\Delta_{n}]$, say, we can bound the integral by
$\ll\frac{1}{r_{n}^{2}}\frac{t}{t-r_{n}}\int_{\Delta_{n}/e^{3}}^{\Delta_{n}}\frac{\mathrm{d}\Delta}{\Delta}\ll\frac{1}{r_{n}^{2}}\int_{r_{n}-1}^{r_{n}+2}\mathrm{d}\log(\Delta^{-1})\ll\frac{1}{r_{n}^{2}}.$
Using the substitution $u=r_{\Delta}$, we split the remaining range into $u\in[r_{n}+2,t/2)$ and $u\in[t/2,t)$. The contribution from the latter is
$\ll\frac{1}{r_{n}^{2}}\int_{t/2}^{t}\frac{t}{(t-r_{\Delta})+1}\frac{\log(u-r_{n})^{5}}{(u-r_{n})^{2}+1}\mathrm{d}u$
$\ll t\log(t)^{5}\int_{t/2}^{t-1}\frac{1}{u^{2}(t-u)}\mathrm{d}u$
$\ll\frac{\log(t)^{6}}{t}\to 0\text{ as }t\to\infty.$
On the former, $t/(t-u)\leq 2$ and we conclude that
$\limsup_{t\to\infty}\int_{e^{-t}}^{\Delta_{n}}\mathcal{F}_{n,t}(\Delta)\mathrm{d}\Delta$ $\ll_{A}\frac{1}{r_{n}^{2}}\int_{r_{n}+2}^{t/2}\frac{\log(u-r_{n})^{4}}{(u-r_{n})^{2}}\mathrm{d}u$
$\ll_{A}\frac{1}{r_{n}^{2}}\Big(1+\int_{1}^{\infty}\frac{\log(x)^{4}}{x^{2}}\mathrm{d}x\Big).$
@@
$\lim_{t\to\infty}\int_{n^{-a}}^{n^{-b}}e^{in\Delta}\mathcal{F}_{n,t}(\Delta)\mathrm{d}\Delta\ll\frac{1}{r_{n}^{2}}\int_{b\log n}^{a\log n}\frac{\log(u-r_{n})^{4}}{(u-r_{n}+1)^{2}}\mathrm{d}u\ll\frac{1}{(\log n)^{2}}$
@@
$\lim_{t\to\infty}\int_{n^{-s_{2}(n)}}^{n^{-s_{1}(n)}}e^{in\Delta}\mathcal{F}_{n,t}(\Delta)\mathrm{d}\Delta=o\bigg(\frac{1}{(\log n)^{2}}\bigg).$
This suggests that the main contribution to $\mathbb{E}[\mathcal{C}_{n,t}\mathrm{(II)}]$ (and in turn $\mathbb{E}[|c_{n,t}|^{2};\mathcal{E}_{n,t}^{(A)}]$) comes from pairs $(\theta_{1},\theta_{2})$ for which $\Delta\approx n^{-a}$ for all scales $a>\delta$.
@@ 5. Upper bound for $\mathcal{F}_{n,t}(\Delta)$
This section proves Proposition 4.3. To begin with, by independence of increments, the expectation in the definition of $\mathcal{F}_{n,t}$ splits into
(5.1) $\mathcal{F}_{n,t}(\Delta)=te^{-2t}\mathbb{E}\Big[e^{\sqrt{2}(X_{r_{n}}(0)+X_{r_{n}}(\Delta))}\mathbf{1}\big\{X_{r_{n}}(*)\leq m(r_{n})+A,*\in\{0,\Delta\}\big\}\Big]$
$\times\mathbb{E}\big[e^{\sqrt{2}(X_{t-r_{n}}^{(r_{n})}(0)+X_{t-r_{n}}^{(r_{n})}(\Delta))}\mathbf{1}\big\{X_{s}^{(r_{n})}(*)\leq U(s)+A,\forall s\in[0,t-r_{n}],*\in\{0,\Delta\}\big\}\Big].$
By Girsanov’s theorem, the first of the two expectations is bounded above by
$e^{2K_{r_{n}}(\Delta)+2r_{n}}\mathbb{Q}_{\Delta}\big(X_{r_{n}}(0)\leq m(r_{n})+A\big),$
where $\mathbb{Q}_{\Delta}$ is a probability measure under which the mean of $X_{r_{n}}(0)$ is shifted by $\sqrt{2}(r_{n}+K_{r_{n}}(\Delta))$ (see Equation 4.9). Since the latter is also Gaussian with variance $r_{n}$, it follows from that the above is
(5.2) $e^{2K_{r_{n}}(\Delta)+2r_{n}}\mathbb{P}\Big(\mathcal{N}\big(0,r_{n}\big)\leq-\sqrt{2}K_{r_{n}}(\Delta)-\frac{3}{2\sqrt{2}}\log(r_{n})+A\Big).$
For $\Delta\leq\Delta_{n}$, $K_{r_{n}}(\Delta)=r_{n}+O(1)$ (cf.  Equation (2.2)) and (5.2) is therefore bounded by
(5.3) $\ll e^{4r_{n}}\int_{-\infty}^{O_{A}(1)-\sqrt{2}r_{n}-\frac{3}{2\sqrt{2}}\log r_{n}}\frac{e^{-u^{2}/(2r_{n})}}{\sqrt{r_{n}}}\mathrm{d}u\ll_{A}\frac{e^{3r_{n}}}{r_{n}^{2}}$
for large enough $n\geq 0$.
We now move on to bounding
(5.4) $\mathbb{E}\big[e^{\sqrt{2}(X_{t-r_{n}}^{(r_{n})}(0)+X_{t-r_{n}}^{(r_{n})}(\Delta))}\mathbf{1}\big\{X_{s}^{(r_{n})}(*)\leq U(s)+A,\forall s\in[0,t-r_{n}],*\in\{0,\Delta\}\big\}\Big]$
beginning with another application of Girsanov’s theorem, by which the above equals
(5.5) $e^{2(t-r_{n})+2(K_{t}(\Delta)-K_{r_{n}}(\Delta))}$
$\times\mathbb{P}$ $\Big(X_{s}^{(r_{n})}(*)\leq A+{U}(s)-2\sqrt{2}s-\sqrt{2}\big(K_{s+r_{n}}(\Delta)-K_{r_{n}}(\Delta)\big),\begin{subarray}{c}\forall s\in[0,t-r_{n}],\\ *\in\{0,\Delta\}\end{subarray}\Big).$
@@
$\ll_{A}\frac{e^{2(t-r_{n})+(r_{\Delta}-r_{n})}\cdot\log(r_{\Delta}-r_{n})^{5}}{(r_{\Delta}-r_{n})^{2}\big((t-r_{\Delta})+1\big)},$
which, combined with (5.3) yields the desired estimate for $\mathcal{F}_{n,t}(\Delta)$, namely
$\mathcal{F}_{n,t}(\Delta)\ll_{A}\frac{1}{r_{n}^{2}}\frac{1}{\Delta}\cdot\frac{\log(r_{\Delta}-r_{n})^{5}}{(r_{\Delta}-r_{n})^{2}}\frac{t}{\big((t-r_{\Delta})+1\big)}.$
To estimate the probability in (5.5), we argue similarly to the proof of Proposition 2.6 in [23] but require a slightly sharper estimate, which we obtain by utilizing the lower order terms in $U(s)$. In particular, we will need the following “linearized” version of $U(s)$.
Recall that
$r_{\Delta}:=t\land\big((\log|\Delta|^{-1})\lor r_{n}\big),\quad\text{$r_{\Delta}\in[r_{n},t]$ for $\Delta\in[0,\Delta_{n}]$}.$
Whenever $r_{\Delta}-r_{n}\geq e$, we define the slope
$\alpha_{\Delta}=\alpha=\bigg(\sqrt{2}-\frac{\log(r_{\Delta}-r_{n})}{2\sqrt{2}(r_{\Delta}-r_{n})}\bigg),$
@@
(5.6) $U(s)\leq\alpha s+\frac{4}{\sqrt{2}}{\log\log(r_{\Delta}-r_{n})}\,\text{ for all $s\in[e,r_{\Delta}-r_{n}]$},$
with equality at $s=r_{\Delta}-r_{n}$.
We are now equipped to bound the probability in (5.5), assuming for now that $(r_{\Delta}-r_{n})\geq e$ so that $U(r_{\Delta}-r_{n})$ is well-defined. Observe first that by symmetry, we can add the restriction $X_{r_{\Delta}-r_{n}}^{(r_{n})}(0)\leq X_{r_{\Delta}-r_{n}}^{(r_{n})}(\Delta)$ to the event in (5.5) since doing so can only change its probability by a factor of $1/2$. Secondly, we can use the estimate
$K_{s+r_{n}}(\Delta)-K_{r_{n}}(\Delta)=s\land(r_{\Delta}-r_{n})+O(1),$
in (5.5), which follows from Equation (2.5). It follows that there exists a constant $C(A)=C$ for which the intersection of $\{X_{r_{\Delta}-r_{n}}^{(r_{n})}(0)\leq X_{r_{\Delta}-r_{n}}^{(r_{n})}(\Delta)\}$ with the events in (5.5) is contained in $E_{1}\cap E_{2}\cap E_{3}$, where
$E_{1}$ $:=\Big\{X_{s}^{(r_{n})}(0)\leq C+\overline{U}(s),\forall s\in[0,r_{\Delta}-r_{n}]\Big\}$
$E_{2}$ $:=\Big\{X_{s}^{(r_{\Delta})}(0)\leq C+\overline{U}(r_{\Delta}-r_{n})-X_{r_{\Delta}-r_{n}}^{(r_{n})}(0),\forall s\in[0,t-r_{\Delta}]\Big\}$
$E_{3}$ $:=\Big\{X_{s}^{(r_{\Delta})}(\Delta)\leq C+\overline{U}(r_{\Delta}-r_{n})-X_{r_{\Delta}-r_{n}}^{(r_{n})}(0),\forall s\in[0,t-r_{\Delta}]\Big\}$
and $\overline{U}(s):=U(s)-2\sqrt{2}s$.
Note that to get the inequality in $E_{3}$, we used (and subsequently dropped) the restriction $X_{r_{\Delta}-r_{n}}^{(r_{n})}(0)\leq X_{r_{\Delta}-r_{n}}^{(r_{\Delta})}(\Delta)$. We have also used the fact that $\overline{U}$ is decreasing on $e,\infty)$.
By Equation (2.10) we have
$\mathbb{P}\Big(E_{2}\cap E_{3}\,\Big|\,X_{r_{\Delta}-r_{n}}^{(r_{n})}(0)=z\Big)\ll\frac{(C+\overline{U}(r_{\Delta}-r_{n})-z+1)^{2}}{((t-r_{\Delta})+1)}.$
Furthermore, by (5.6),
$\mathbb{P}\Big(E_{1}\Big|X_{r_{\Delta}-r_{n}}^{(r_{n})}(0)=z\Big)$
$\leq\mathbb{P}\bigg(X_{s}^{(r_{n})}(0)\leq C+\frac{4}{\sqrt{2}}{\log\log(r_{\Delta}-r_{n})}+(\alpha-2\sqrt{2})s,$
$\hskip 200.74998pt\forall s\in[e,r_{\Delta}-r_{n}]\Big|X_{r_{\Delta}-r_{n}}^{(r_{n})}(0)=z\bigg),$
@@
and we can estimate the right-hand side using Equation (2.11). Indeed, it is bounded by the probability that
$Z_{s}:=C+\frac{4}{\sqrt{2}}\log\log(r_{\Delta}-r_{n})+\bigg(\alpha-2\sqrt{2}-\frac{z}{(r_{\Delta}-r_{n})}\bigg)s-X_{s}^{(r_{n})}(0)+\frac{z}{(r_{\Delta}-r_{n})}s,$
which is a Brownian bridge from $C+(4/\sqrt{2})\log\log(r_{\Delta}-r_{n})$ to $(C+\overline{U}(r_{\Delta}-r_{n})-z)$, remains positive at all times $s\in[e,r_{\Delta}-r_{n}]$. This gives
$\ll\frac{(C+\log(r_{\Delta}-r_{n}))(C+\overline{U}(r_{\Delta}-r_{n})-z+1)}{((r_{\Delta}-r_{n})+1)},$
where we bounded the iterated logarithm by $\log(r_{\Delta}-r_{n})$ for simplicity.
By integrating over the value of $C+\overline{U}(r_{\Delta}-r_{n})-X_{r_{\Delta}-r_{n}}^{(r_{n})}(0)$, we can then bound the quantity in (5.5) by
$e^{2(t-r_{n})+2(K_{t}(\Delta)-K_{r_{n}}(\Delta))}\mathbb{E}\bigg[\mathbf{1}_{E_{1}}\mathbb{P}\Big(E_{2}\cap E_{3}\,\Big|\,X_{r_{\Delta}-r_{n}}^{(r_{n})}(0)\Big)\bigg]$
$\ll e^{2(t-r_{n})+2(r_{\Delta}-r_{n})}$
@@
$\times\int_{0}^{\infty}\frac{e^{-\frac{(x-C-\overline{U}(r_{\Delta}-r_{n}))^{2}}{2(r_{\Delta}-r_{n})}}}{\sqrt{r_{\Delta}-r_{n}}}\frac{(C+\log(r_{\Delta}-r_{n}))(x+1)(x+1)^{2}}{\big((r_{\Delta}-r_{n})+1\big)\big((t-r_{\Delta})+1\big)}\mathrm{d}x$
(5.7) $\ll_{A}e^{2(t-r_{n})+2(r_{\Delta}-r_{n})}\frac{\log(r_{\Delta}-r_{n})e^{-(\overline{U}(r_{\Delta}-r_{n})+C)^{2}/2(r_{\Delta}-r_{n})}}{(r_{\Delta}-r_{n})^{3/2}\big((t-r_{\Delta})+1\big)}.$

### v2

@@ 4.1. Initial reductions
It therefore suffices to prove that the right-hand side is $\ll(\log n)^{-1/2+\varepsilon}$ as $t\to\infty$. The bulk of the work lies in proving the following uniform upper bound for $\mathcal{F}_{n,t}$ whose proof is deferred to Section 5.
To state it, we introduce the branching time
${r(\Delta)}=t\land\big(\log|\Delta|^{-1}\big).$
@@
Let $\mathcal{F}_{n,t}$ be defined as in Equation (4.23). Then there exists a constant $C(A)=C>0$ such that for large enough $t,n>0$, $\mathcal{F}_{n,t}(\Delta)$ satisfies the bound
(4.25) $\mathcal{F}_{n,t}(\Delta)\leq C\sqrt{r_{n}}\cdot\frac{e^{r(\Delta)}}{r(\Delta)}\cdot\frac{t}{(t-{r(\Delta)})+1}\cdot\frac{\log\big(r(\Delta)\big)^{4}}{(r(\Delta)-r_{n})\lor 1}$
uniformly in $\Delta\in[0,\Delta_{n}]$.
@@
For large enough $n>0$, we can use Proposition 4.3 to bound the integral in (4.24), beginning with the range $0,e^{-t}]$ where
$\limsup_{t\to\infty}\int_{0}^{e^{-t}}\mathcal{F}_{n,t}(\Delta)\mathrm{d}t\ll_{A}\sqrt{r_{n}}\lim_{t\to\infty}te^{t-t}\frac{1}{t}\bigg(\frac{\log(t)^{4}}{t-r_{n}}\bigg)=0.$
On $e^{-t},\Delta_{n}]$, Proposition 4.3 yields
$\int_{e^{-t}}^{\Delta_{n}}\mathcal{F}_{n,t}(\Delta)\mathrm{d}\Delta\ll_{A}\sqrt{r_{n}}\int_{e^{-t}}^{\Delta_{n}}\frac{1}{r(\Delta)}\frac{t}{(t-{r(\Delta)})+1}\frac{\log(r(\Delta))^{4}}{(r(\Delta)-r_{n})\lor 1}\frac{1}{\Delta}\mathrm{d}\Delta.$
We first take care of the range where $r(\Delta)$ is close to $r_{n}$. On $\Delta_{n}/e^{3},\Delta_{n}]$, say, we have
$\limsup_{t\to\infty}\int_{\Delta_{n}/e^{3}}^{\Delta_{n}}\mathcal{F}_{n,t}(\Delta)\mathrm{d}\Delta$
$\ll\limsup_{t\to\infty}\frac{\sqrt{r_{n}}\log(r_{n}+2)^{4}}{r_{n}-1}\frac{t}{t-r_{n}+2}\int_{\Delta_{n}/e^{3}}^{\Delta_{n}}\frac{\mathrm{d}\Delta}{\Delta}$
$\ll\frac{(\log r_{n})^{4}}{\sqrt{r_{n}}}\int_{r_{n}-1}^{r_{n}+2}\mathrm{d}\log(\Delta^{-1})\ll\frac{(\log r_{n})^{4}}{\sqrt{r_{n}}}.$
Using the substitution $u=r(\Delta)=\log\Delta^{-1}$, we split the remaining range into $u\in[r_{n}+2,t/2)$ and $u\in[t/2,t)$. The contribution from the latter is
$\ll\sqrt{r_{n}}\int_{t/2}^{t}\frac{1}{u}\frac{t}{(t-u)+1}\frac{\log(u)^{4}}{(u-r_{n})\lor 1}\mathrm{d}u$
$\ll\sqrt{r_{n}}\frac{t}{t-1}\frac{\log(t)^{4}}{(t-1-r_{n})}+\sqrt{r_{n}}t\log(t)^{4}\int_{t/2}^{t-1}\frac{1}{u(u-r_{n})(t-u)}\mathrm{d}u$
$\ll\sqrt{r_{n}}\frac{\log(t)^{5}}{t}\to 0\text{ as }t\to\infty.$
On the former, we have $t/(t-u)\leq 2$ and we conclude that
$\limsup_{t\to\infty}\int_{e^{-t}}^{\Delta_{n}}\mathcal{F}_{n,t}(\Delta)\mathrm{d}\Delta$ $\ll_{A}\sqrt{r_{n}}\int_{r_{n}+2}^{\infty}\frac{1}{u}\frac{\log(u)^{4}}{(u-r_{n})}\mathrm{d}u$
$\ll_{A}\frac{\log(2r_{n})^{4}}{\sqrt{r_{n}}}\int_{2}^{r_{n}}\frac{\mathrm{d}x}{x}+\sqrt{r_{n}}\int_{r_{n}}^{\infty}\frac{\log(2x)^{4}}{x^{2}}\mathrm{d}x$
$\ll_{A}\frac{(\log\log n)^{5}}{\sqrt{\log n}},$
which is $\ll_{A}(\log n)^{-1/2+\varepsilon}$ for any $\varepsilon>0$ as claimed. ∎
@@
$\lim_{t\to\infty}\int_{n^{-a}}^{n^{-b}}e^{in\Delta}\mathcal{F}_{n,t}(\Delta)\mathrm{d}\Delta\ll_{A}\frac{1}{(\log n)^{1/2-\varepsilon}}$
@@
$\lim_{t\to\infty}\int_{n^{-s_{2}(n)}}^{n^{-s_{1}(n)}}e^{in\Delta}\mathcal{F}_{n,t}(\Delta)\mathrm{d}\Delta=o_{A}\bigg(\frac{1}{(\log n)^{1/2-\varepsilon}}\bigg).$
This suggests that the main contribution to $\mathbb{E}[\mathcal{C}_{n,t}\mathrm{(II)}]$ (and in turn $\mathbb{E}[|c_{n,t}|^{2};{E}_{n,t}{(A)}]$) comes from pairs $(\theta_{1},\theta_{2})$ for which $|\theta_{2}-\theta_{1}|\approx n^{-a}$ for all scales $a>\delta$.
@@ 5. Upper bound for $\mathcal{F}_{n,t}(\Delta)$
This section proves Proposition 4.3. Recalling that
$\mathcal{F}_{n,t}(\Delta)=te^{-2t}\mathbb{E}\big[e^{\sqrt{2}(X_{t}(0)+X_{t}(\Delta))}\mathbf{1}\big\{X_{s}(*)\leq U(s)+A,\forall s\in[r_{n},t],*\in\{0,\Delta\}\big\}\Big]$
an application of Girsanov’s theorem shows that $\mathcal{F}_{n,t}(\Delta)$ equals
(5.1) $te^{2K_{t}(\Delta)}\mathbb{P}\Big(X_{s}(*)\leq A+{U}(s)-\sqrt{2}(s+K_{s}(\Delta)),{\forall s\in[r_{n},t],*\in\{0,\Delta\}}\Big).$
@@
$\ll_{A}\sqrt{r_{n}}\frac{e^{r(\Delta)}}{r(\Delta)}\frac{t}{(t-r(\Delta))+1}\cdot\frac{\log(r(\Delta)-r_{n})^{4}}{(r(\Delta)-r_{n})\lor 1},$
by arguing similarly to the proof of Proposition 2.6 in [22]. We require a slightly sharper estimate than the one proved there, which we obtain by utilizing the lower order terms in $U(s)$.
Assume for now that $r(\Delta)-r_{n}>1$. Observe first that by symmetry, we can add the restriction $X_{r(\Delta)}(0)\leq X_{r(\Delta)}(\Delta)$ to the event in (5.1) since doing so can only change its probability by a factor of $1/2$. Secondly, we can use the estimate
$K_{s}(\Delta)=s\land(\log|\Delta|^{-1})+O(\|k^{\prime}\|_{\infty})=s\land r(\Delta)+O(1),$
(cf. Equation (2.2)), which is uniform in $s\geq 0$ and $\Delta\in[0,\Delta_{n}]$. It follows that there exists a constant $C=C(A)>0$ for which the intersection of $\{X_{r(\Delta)}(0)\leq X_{r(\Delta)}(\Delta)\}$ with the event in (5.1) is contained in $E_{1}\cap E_{2}\cap E_{3}$, where
$E_{1}$ $:=\Big\{X_{s}(0)\leq C+U(s)-2\sqrt{2}s,\forall s\in[r_{n},r(\Delta)]\Big\}$
$E_{2}$ $:=\Big\{X_{s}^{(r(\Delta))}(0)\leq C-\sqrt{2}r(\Delta)-X_{r(\Delta)}(0),\forall s\in[0,t-r(\Delta)]\Big\}$
$E_{3}$ $:=\Big\{X_{s}^{(r(\Delta))}(\Delta)\leq C-\sqrt{2}r(\Delta)-X_{r(\Delta)}(0),\forall s\in[0,t-r(\Delta)]\Big\}$
Note that to get the inequality in $E_{3}$, we used (and subsequently dropped) the restriction $X_{r(\Delta)}(0)\leq X_{r(\Delta)}(\Delta)$.
We thus need to bound
(5.2) $\mathbb{P}(E_{1}\cap E_{2}\cap E_{3})=\mathbb{E}\Big[\mathbf{1}_{E_{1}}\mathbb{P}\big(E_{2}\cap E_{3}\big|\Sigma_{\Delta}\big)\Big],$
where $\Sigma_{\Delta}=\sigma((X_{s}(\theta))_{\theta\in[0,1],s\leq r(\Delta)})$. This will require a “linearized” version of U(s){U}(s). For $r(\Delta)\geq e$, define the slope
$\alpha_{\Delta}=\alpha=\bigg(\sqrt{2}-\frac{\log r(\Delta)}{2\sqrt{2}r(\Delta)}\bigg),$
@@
(5.3) ${U}(s)\leq\alpha s+\frac{4}{\sqrt{2}}{\log\log r(\Delta)}\,\text{ for all $s\in[e,r(\Delta)]$},$
with equality at $s=r(\Delta)$. For simplicity, we also define
$V_{\Delta,A}(s)=V(s)=C+\frac{4}{\sqrt{2}}\log\log r(\Delta)+(\alpha-2\sqrt{2})s$
which is greater than $U(s)-2\sqrt{2}s$ for all $s\in[e,r(\Delta)]$.
We can now estimate (5.2). Letting $(B_{s})_{s\leq r(\Delta)}$ denote a standard Brownian motion, we note on the one hand that
$\mathbb{P}\Big(E_{1}\,\Big|\,X_{r_{n}}(0)=y,X_{r(\Delta)}(0)=z\Big)$
$\leq\mathbb{P}\Big(X_{s-r_{n}}^{(r_{n})}(0)\leq V(s),\forall s\in[0,r(\Delta)-r_{n}]\,\Big|\,X_{r_{n}}(0)=y,X_{r(\Delta)}(0)=z\Big)$
$=\mathbb{P}\big(B_{s}\leq V(s),\forall s\in[0,r(\Delta)-r_{n}]\,\big|\,B_{r_{n}}=y,B_{r(\Delta)}=z\big),$
@@
and that the right-hand side can be estimated using Equation (2.8). Indeed, it is equal to the probability that a Brownian bridge from $(V(r_{n})-y)$ to $(V(r(\Delta))-z)$ remains positive, which is
$\ll\bigg(\frac{(V(r_{n})-y+1)(V(r(\Delta))-z+1)}{(r(\Delta)-r_{n})+1}\land 1\bigg).$
On the other hand, Equation (2.7) yields
$\mathbb{P}\big(E_{2}\cap E_{3}\,\big|\,\Sigma_{\Delta}\big)\ll_{A}\bigg(\frac{(V(r(\Delta))-X_{r(\Delta)}(0)+1)^{2}}{((t-r(\Delta))+1)}\land 1\bigg).$
By integrating over the values of
$(y,x)=(X_{r_{n}}(0),X_{r(\Delta)}(0)-X_{r_{n}}(0))$
we can therefore conclude that $\mathcal{F}_{n,t}(\Delta)$ is
$\ll te^{2K_{t}(\Delta)}\mathbb{E}\Big[\mathbf{1}_{E_{1}}\mathbb{P}\big(E_{2}\cap E_{3}\,\big|\,X_{r(\Delta)}(0)\big)\Big]$
(5.4) $\ll_{A}\frac{te^{2r(\Delta)}}{(r(\Delta)-r_{n})(t-r(\Delta)+1)}$
$\times\int_{-\infty}^{V(r_{n})}\frac{e^{-\frac{y^{2}}{2r_{n}}}}{\sqrt{r_{n}}}\int_{-\infty}^{V(r(\Delta))-y}\frac{e^{-\frac{x^{2}}{2(r(\Delta)-r_{n})}}}{\sqrt{r(\Delta)-r_{n}}}F(y,(x+y))\mathrm{d}{x}\mathrm{d}{y},$
where
$F(a,b):={(V(r_{n})-a+1)}{(V(r(\Delta))-b+1)^{3}}.$
The change of variables $z=x+y$ yields
$\frac{te^{2r(\Delta)}}{(r(\Delta)-r_{n})(t-r(\Delta)+1)}$
$\times\int_{-\infty}^{V(r_{n})}\int_{-\infty}^{V(r(\Delta))}\frac{e^{-z^{2}/2r(\Delta)}}{\sqrt{r(\Delta)}}\frac{e^{-(y-\frac{r_{n}}{r(\Delta)}z)^{2}/2(\tfrac{r_{n}}{r(\Delta)}(r(\Delta)-r_{n}))}}{\sqrt{(r_{n}/r(\Delta))(r(\Delta)-r_{n})}}F(y,z)\mathrm{d}z\mathrm{d}y.$
We first evaluate the integral over $y$, which equals
$\mathbb{E}\Big[\mathbf{1}\big(X<V(r_{n})\big)\big(V(r_{n})-X+1\big)\Big],\text{ where }X\sim\mathcal{N}\Big(\mu=\frac{r_{n}}{r(\Delta)}z,\sigma^{2}=\frac{r_{n}(r(\Delta)-r_{n})}{r(\Delta)}\Big).$
This can be decomposed into
(5.5) $\big(V(r_{n})-\mu+1\big)\mathbb{P}(X<V(r_{n}))+\mathbb{E}\big[\mathbf{1}(X<V(r_{n}))\big(X-\mu\big)\big]$
(5.6) $\leq\big(V(r_{n})-\mu+1\big)+\sigma\mathbb{E}\Big[Z\mathbf{1}\big(Z<(V(r_{n})-\mu)/\sigma\big)\Big]$
(for $Z$ a standard Gaussian), which is equal to
$(V(r_{n})-\mu+1)+\frac{\sigma}{\sqrt{2\pi}}e^{-(V(r_{n})-\mu)^{2}/(2\sigma^{2})}.$
Inserting this into (5.5) yields the bound
$\ll$ $\frac{te^{2r(\Delta)}}{(r(\Delta)-r_{n})(t-r(\Delta)+1)}$
$\times\bigg(\int_{-\infty}^{V(r(\Delta))}\frac{e^{-z^{2}/2r(\Delta)}}{\sqrt{r(\Delta)}}\Big(V(r_{n})-\frac{r_{n}}{r(\Delta)}z+1\Big)\big(V(r(\Delta))-z+1\big)^{3}\mathrm{d}z$
@@
$\hskip 20.075pt+\sigma\int_{-\infty}^{V(r(\Delta))}\frac{e^{-z^{2}/2r(\Delta)-(V(r_{n})-\frac{r_{n}}{r(\Delta)}z)^{2}/2\sigma^{2}}}{\sqrt{r(\Delta)}}\big(V(r(\Delta))-z+1\big)^{3}\mathrm{d}z\bigg)$
which by the substitution $\bar{z}:=V(r(\Delta))-z$ equals
$\frac{te^{2r(\Delta)}}{(r(\Delta)-r_{n})(t-r(\Delta)+1)}\bigg(\int_{0}^{\infty}\frac{e^{-(V(r(\Delta))-\bar{z})^{2}/2r(\Delta)}}{\sqrt{r(\Delta)}}\Big(\frac{r_{n}}{r(\Delta)}\bar{z}+1\Big)\big(\bar{z}+1\big)^{3}\mathrm{d}\bar{z}$
$\hskip 50.18748pt+\sigma\int_{0}^{\infty}\frac{e^{-(V(r(\Delta))-\bar{z})^{2}/2r(\Delta)-\bar{z}^{2}/2(r_{\Delta}-r_{n})}}{\sqrt{r(\Delta)}}\big(\bar{z}+1\big)^{3}\mathrm{d}\bar{z}\bigg)$
$\leq\frac{(1+\sigma)te^{2r(\Delta)}}{(r(\Delta)-r_{n})(t-r(\Delta)+1)}\int_{0}^{\infty}\frac{e^{-(V(r(\Delta))-\bar{z})^{2}/2r(\Delta)}}{\sqrt{r(\Delta)}}(\bar{z}+1)^{4}\mathrm{d}\bar{z}$
(5.7) $\leq\frac{(1+\sigma)te^{2r(\Delta)}}{(r(\Delta)-r_{n})(t-r(\Delta)+1)}\frac{e^{-V(r(\Delta))^{2}/2r(\Delta)}}{\sqrt{r(\Delta)}}\int_{0}^{\infty}e^{\tfrac{V(r(\Delta))}{r(\Delta)}\bar{z}}(\bar{z}+1)^{4}\mathrm{d}\bar{z}.$